Find the velocity aquired by an electron in falling through a potential difference of 2000 volts.


1
Expert's answer
2022-06-06T08:53:43-0400
"\frac{mv^2}{2}=eV"

"v=\sqrt{\frac{2eV}{m}}=\sqrt{\frac{2*1.6*10^{-19}*2000}{9.1*10^{-31}}}=2.65*10^7\:\rm m/s"